# FabGPT: An Efficient Large Multimodal Model for Complex Wafer Defect Knowledge Queries

### v1

@@ Abstract.
Intelligence is key to advancing integrated circuit (IC) fabrication. Recent breakthroughs in Large Multimodal Models (LMMs) have unlocked unparalleled abilities in understanding images and text, fostering intelligent fabrication. Leveraging the power of LMMs, we introduce FabGPT, a customized IC fabrication large multimodal model for wafer defect knowledge query. FabGPT manifests expertise in conducting defect detection in Scanning Electron Microscope (SEM) images, performing root cause analysis, and providing expert question-answering (Q&A) on fabrication processes. FabGPT matches enhanced multimodal features to automatically detect minute defects under complex wafer backgrounds and reduce the subjectivity of manual threshold settings. Besides, the proposed modulation module and interactive corpus training strategy embed wafer defect knowledge into the pre-trained model, effectively balancing Q&A queries related to defect knowledge and original knowledge and mitigating the modality bias issues. Experiments on in-house fab data (SEM-WaD) show that our FabGPT achieves significant performance improvement in wafer defect detection and knowledge querying.
# Corresponding authors: {qisunchn, czhuo}@zju.edu.cn
@@ 1. Introduction
To address the issues mentioned above, we propose an efficient LMM, FabGPT, that employs a three-stage strategy: modal enhancement, detection, and Q&A stage. This strategy allows us to build on the inherent capabilities of pre-trained models by embedding high-quality prompt instructions, enabling it to detect wafer defects in the IC domain and query the related knowledge. Additionally, to further alleviate “modality bias”, we introduce an interactive corpus training strategy.
Our contributions are summarized as follows:
@@ 2.1. IC Wafer Defect Analysis
2.2. Collection and Processing of Dataset
Due to the absence of publicly available datasets for wafer surface defects, we collect an in-house dataset (SEM-WaD) using SEM techniques from various IC manufacturing steps and products. It comprises 1,226 defect-free images and 1,182 images with four common types of defects (holes: 250, particles: 500, pattern deformities: 250, scratches: 182) from our fab partners. The wafer images in SEM-WaD contain diverse and complex backgrounds, with each defect type exhibiting unique morphologies and features. Consequently, we meticulously annotate each image in the dataset with details such as image IDs and production steps. Some examples are shown in Figure 2.
@@ 2.4. Fine-Tuning Methods
Fine-tuning methods primarily include full (alaparthi2020bidirectional, ; yuan2021tokens, ) and partial fine-tuning (rebuffi2017learning, ; lai2023lisa, ; gu2023anomalygpt, ), which involves retraining parts of a model pre-trained on large datasets to adapt to specific tasks. In full fine-tuning, all pre-trained model layer weights are trainable, enabling flexible adaptation to new data features by adjusting all parameters. For example, BERT (alaparthi2020bidirectional, ) is trained on large corpora adjusting all parameters in its bidirectional Transformer structure to optimize tasks such as sentiment analysis question answering and text summarization. ViT (yuan2021tokens, ) adjusts its Transformer structure by specialized visual datasets, efficiently adapting to tasks such as image classification and object detection. Full fine-tuning enhances model adaptation to specific downstream tasks but also poses a higher risk of overfitting and increasing computational costs, limiting its wide applications.
In partial fine-tuning (rebuffi2017learning, ; lai2023lisa, ; gu2023anomalygpt, ), it adjusts only partial parameters in pre-trained models and keeps the rest fixed. Recent popular partial fine-tuning strategies include Adapters, Prompt Learning, and LoRA. For example, Rebuffi et al. (rebuffi2017learning, ) introduced residual adapters to neural networks, optimizing only these modules to reduce fine-tuning costs due to their fewer parameters. Lai et al. (lai2023lisa, ) applied the LoRA method to consistently update the bottom embeddings and top linear head of the pre-trained model, while randomly updating a few intermediate self-attention layers to understand input text for precise segmentation. Gu et al. proposed AnomalyGPT (gu2023anomalygpt, ), which fine-tunes an LLM using embedded prompt instructions to identify types and locations of defects.
@@ 3.2. Modal Enhancement Stage
Figure 4. The Prediction Module (PM).
@@
The pre-trained encoder captures pixel-level detail features in the latent space, but its repeated down-sampling operations result in the loss of semantic features. As shown in Figure 4, we design the PM to predict defect categories in the image and enhance semantic features by embedding the expected result $p_{n}$ into the initial vectors.
@@ 3.4. Q&A Stage
The quality of prompt instructions significantly affects the model’s ability to understand knowledge and respond to queries. Thus, it is necessary to align enhanced multimodal features to capture visual information and alleviate the model’s fine-tuning burden. Inspired by Q-Former (li2023blip, ), a bidirectional self-attention in Figure 5 allows $T_{img}$ to absorb semantic and detailed information from $T_{mas}$ to obtain $f_{i\sim m}$, facilitating interaction within the same modality, formulated as:
@@ 3.5. Loss Functions
We employ three loss functions to constrain the detection and dialogue processes. Focal Loss (lin2017focal, ) and Dice Loss (milletari2016v, ) are used to improve the model’s segmentation and localization abilities, and Cross-Entropy Loss is used to improve the model’s classification and Q&A ones.
@@ 4.1. Experimental Setups
Table 1. The results of the Image-AUC and Pixel-AUC metrics from fully supervised experiments on the SEM-WaD dataset. The best and second-best methods are highlighted in red and blue fonts, respectively.
Category DevNet (pang2021explainable, ) DRA (ding2022catching, ) BGAD (yao2023explicit, ) PRN (zhang2023prototypical, ) Lisa (lai2023lisa, ) AnomalyGPT (gu2023anomalygpt, ) FabGPT (ours)
Hole 63.92 / 80.10 81.44 / 85.75 35.78 / 87.10 79.26 / 84.30 87.37 / 84.31 91.37 / 93.68 94.28 / 97.03
Particle 62.20 / 86.30 96.71 / 96.86 81.87 / 92.75 80.41 / 87.60 93.34 / 91.73 92.64 / 94.39 94.43 / 97.30
Scratch 19.84 / 76.74 89.94 / 75.28 65.27 / 88.98 76.93 / 77.02 84.45 / 85.28 89.58 / 92.58 90.32 / 95.70
Pattern Deformation 65.02 / 48.78 85.52 / 90.71 72.37 / 81.90 75.87 / 76.53 85.51 / 87.72 86.23 / 89.66 88.19 / 92.40
Average 52.74 / 72.98 88.14 / 87.15 63.82 / 87.68 78.12 / 81.36 87.67 / 87.26 89.96 / 92.58 91.81 / 95.61
Table 2. The results of the PRO and AP metrics from fully supervised experiments on the SEM-WaD dataset. The best and second-best methods are highlighted in red and blue fonts, respectively.
Category DevNet (pang2021explainable, ) DRA (ding2022catching, ) BGAD (yao2023explicit, ) PRN (zhang2023prototypical, ) Lisa (lai2023lisa, ) AnomalyGPT (gu2023anomalygpt, ) FabGPT (ours)
Hole - / 86.98 57.81 / - 80.81 / 51.33 80.90 / 76.48 88.66 / 79.18 86.39 / 83.47 90.01 / 85.69
Particle - / 92.38 66.67 / - 89.58 / 51.33 88.84 / 84.79 83.57 / 72.87 89.74 / 86.02 92.28 / 91.58
Scratch - / 50.55 25.83 / - 75.33 / 52.28 67.82 / 51.70 57.77 / 35.73 80.32 / 60.32 83.00 / 79.19
Pattern Deformation - / 82.07 60.71 / - 81.08 / 54.75 74.63 / 70.57 77.75/ 69.01 84.35 / 80.73 87.40 / 86.72
Average - / 77.80 52.76 / - 81.70 / 52.66 78.05 / 70.89 76.94/ 64.20 85.20 / 77.64 88.17 / 85.80
@@
Our experiment is conducted on the in-house SEM-WaD dataset. The dataset comprises images with a resolution of $480\times 480$, each accompanied by a corresponding mask image and related textual descriptions and analyses. We divide the data into training and test sets in a 7:3 ratio, where both sets consist of good and defective images.
@@
These baselines are also tuned on our SEM-WaD dataset, following their publicly available implementations and models. Among these baselines, Lisa and AnomalyGPT are LMMs that support Q&A while the others do not.
@@ 4.2. Quantitative Results
The quantitative results report the accuracy of the supervised defect detection and the correctness of the knowledge Q&A responses in detail.
@@
Table 1 and Table 2 report the Image-AUC, Pixel-AUC, PRO, and AP values for different methods across 4 categories within the SEM-WaD dataset. It can be observed that our model outperforms all other methods in four evaluation metrics for most defect categories. For example, compared to AnomalyGPT (gu2023anomalygpt, ), which also employs prompt learning for fine-tuning, our model achieves a higher Image-AUC by 1.85% and a higher Pixel-AUC by 3.03%. Compared to the traditional detection model PRN (zhang2023prototypical, ), our model surpasses it by 10.12% in PRO and 14.91% in AP.
Table 3. The accuracy (%) in the models’ responses to different questions. The best and second-best methods are highlighted in red and blue fonts, respectively. “-” denotes the question is unsupported.
Methods Defect-Related Unrelated All
Presence Category Location Quantity Description Analysis
Lisa (lai2023lisa, ) 95.00 92.50 - - 72.50 - 20.00 70.00
AnomalyGPT (gu2023anomalygpt, ) 95.00 75.00 72.50 77.50 82.50 - 12.00 69.08
GPT-4 (achiam2023gpt, ) 100.00 37.50 87.50 80.00 85.00 90.00 98.00 82.57
FabGPT (ours) 100.00 97.50 95.00 95.00 95.00 97.50 98.00 96.86
Figure 6. Visualization results of four defect types, compared with non-LMM baseline models.
@@ 4.3. Qualitative Results
We produce comparison figures for detection results and dialogue diagrams for Q&A results on the SEM-WaD dataset. Intuitive detection results are validated using heatmaps and mask images to demonstrate the effectiveness of our model in the supervised detection task.
@@
According to Figure 8 and Figure 9, the proposed FabGPT offers detailed descriptions of defect knowledge, providing engineers with more useful information. It should be noted that the questions shown in the figures are not included in our training Corpus-A and Corpus-B.
Specifically, FabGPT is proficient in identifying the defect in SEM images, localizing its position, and discerning specific properties, such as “hole” evident in Figure 8 and “scratch” depicted in Figure 9. In addition, FabGPT can analyze the root causes of wafer defects and propose appropriate solutions from a process perspective, thereby serving as an effective compass for refining the production line.
Furthermore, FabGPT demonstrates competency in providing accurate answers to general IC questions that are not closely related to the input images, depicted in Figure 9, such as explaining the process steps in IC manufacturing, the steps of lithography, and how AI is advancing the development of IC manufacturing. This verifies the effectiveness of our proposed modulation module in resolving modality bias.
@@ 4.4. Ablation Studies
Figure 7. Visualization results of four defect types, compared with LMM-based baseline models.
Figure 8. Dialogue example of FabGPT for “hole” type defects.
Figure 9. Dialogue example of FabGPT for “scratch” type defects. Questions with $\ast$” are general IC questions that are not closely related to the input images.
@@
Table 5 shows the advantages of the operation strategy we designed for the PM. It can be seen that compared to matrix multiplication or other operations, using the cosine of the angle to calculate the similarity between two vectors provides a better assessment of their relationship. Moreover, when combined with Linear and ReLU, it achieves a better computational effect.
@@
From Table 6, it can be seen that our embedding scheme for prompt instructions showcases optimal effectiveness in mitigating modality bias issues. By aligning and updating the corresponding coefficient “$a$” for visual tokens, our model effectively discerns the relationship between user queries and visual inputs and mitigates the modality bias issues.
@@ 5. Conclusions
In this paper, we introduce a novel large multimodal language model, FabGPT, for defect knowledge querying in the IC field, including defection, analysis, Q&A, etc. It employs three stages to gradually achieve the functionality of defect detection and high-quality dialogue. Enhanced feature tokens aid the model in automatically conducting high-precision detection. Dynamically aligned and corrected tokens fine-tune the LLM, enabling not only attribution analysis of defect regions and correct responses regarding defect type, location, quantity, etc., but also mitigate the modality bias issues within conversations. We validate the effectiveness of FabGPT on the SEM-WaD dataset and 100 questions. Our work provides great convenience for the semiconductor industry and also offers insights for further LMM research.

### v2

@@ Abstract.
Intelligence is key to advancing integrated circuit (IC) fabrication. Recent breakthroughs in Large Multimodal Models (LMMs) have unlocked extraditionary abilities in understanding images and text, fostering intelligent fabrication. Leveraging the power of LMMs, we introduce FabGPT, a customized IC fabrication large multimodal model for wafer defect knowledge query. FabGPT manifests expertise in conducting defect detection in Scanning Electron Microscope (SEM) images, performing root cause analysis, and providing expert Q&A on fabrication processes. FabGPT matches enhanced multimodal features to automatically detect minute defects under complex wafer backgrounds and reduce the subjectivity of manual threshold settings. Besides, the proposed modulation module and interactive corpus training strategy embed wafer defect knowledge into the pre-trained model, effectively balancing Q&A queries related to defect knowledge and original knowledge and mitigating the modality bias issues. Experiments on in-house fab data show that FabGPT achieves significant performance improvement in wafer defect detection and knowledge querying.
# Corresponding authors
@@ 1. Introduction
To address the issues mentioned above, we propose an efficient LMM, FabGPT, that employs a three-stage strategy: modal enhancement, detection, and Q&A stage. This strategy allows us to build on the inherent capabilities of pre-trained models by embedding high-quality prompt instructions, enabling it to detect wafer defects in the IC domain and query the related knowledge. Additionally, to further alleviate “modality bias”, we introduce an interactive corpus training strategy. Our contributions are summarized as follows:
@@ 2.1. IC Wafer Defect Analysis
We use an in-house dataset, SEM-WaD, which comprises 1,226 defect-free images and 1,182 images with four common types of defects (holes, particles, pattern deformities, and scratches). The wafer images contain diverse and complex backgrounds, with each defect type exhibiting unique morphologies and features. Some examples are shown in Figure 2.
@@ 2.3. Fine-Tuning Methods
Fine-tuning methods primarily include full (alaparthi2020bidirectional, ; yuan2021tokens, ) and partial fine-tuning (rebuffi2017learning, ; lai2023lisa, ; gu2023anomalygpt, ). In full fine-tuning, all pre-trained model layer weights are trainable, enabling flexible adaptation to new data features by adjusting all parameters. For example, BERT (alaparthi2020bidirectional, ) is trained on large corpora adjusting all parameters in its bidirectional Transformer structure to optimize tasks such as sentiment analysis question answering and text summarization. Full fine-tuning improves task adaptation but increases overfitting risk and costs, limiting its use.
In partial fine-tuning (rebuffi2017learning, ; lai2023lisa, ; gu2023anomalygpt, ), it adjusts only partial parameters in pre-trained models and keeps the rest fixed. Recent popular partial fine-tuning strategies include Adapters, Prompt Learning, and LoRA. For example, Lai et al. (lai2023lisa, ) applied the LoRA method to consistently update the bottom embeddings and top linear head of the pre-trained model, while randomly updating a few intermediate self-attention layers to understand input text for precise segmentation. Gu et al. proposed AnomalyGPT (gu2023anomalygpt, ), which fine-tunes an LLM using embedded prompt instructions to identify types and locations of defects.
@@ 3.2. Modal Enhancement Stage
Figure 4. (a) The Prediction Module (PM), (b) The Modulation Module.
@@
The pre-trained encoder captures pixel-level detail features in the latent space, but its repeated down-sampling operations result in the loss of semantic features. As shown in Figure 4 (a), we design the PM to predict defect categories in the image and enhance semantic features by embedding the expected result $p_{n}$ into the initial vectors.
@@ 3.4. Q&A Stage
The quality of prompts impacts the model’s understanding and responses, so aligning multimodal features is needed to capture comprehensive information and reduce the fine-tuning burden. Inspired by Q-Former (li2023blip, ), a bidirectional self-attention in Figure 4 (b) allows $T_{img}$ to absorb semantic and detailed information from $T_{mas}$ to obtain $f_{i\sim m}$, facilitating interaction within the same modality, formulated as:
@@ 3.5. Loss Functions
We employ three loss functions to constrain detection and dialogue processes. Focal Loss (lin2017focal, ) and Dice Loss (milletari2016v, ) are used to improve the model’s segmentation and localization abilities, and Cross-Entropy Loss is used to improve the model’s classification and Q&A ones.
@@ 4.1. Experimental Setups
Table 1. The fully-supervised experimental results. The best and second-best methods are highlighted in red and blue.
Methods Image-AUC / Pixel-AUC PRO / AP
Hole Particle Scratch Pattern Average Hole Particle Scratch Pattern Average
Deformation Deformation
DevNet (pang2021explainable, ) 63.92 / 80.10 62.20 / 86.30 19.84 / 76.74 65.02 / 48.78 52.74 / 72.98     -    / 86.98     -    / 92.38     -    / 50.55     -    / 82.07     -    / 77.80
DRA (ding2022catching, ) 81.44 / 85.75 96.71 / 96.86 89.94 / 75.28 85.52 / 90.71 88.14 / 87.15 57.81 /     - 66.67 /     - 25.83 /     - 60.71 /     - 52.76 /     -
BGAD (yao2023explicit, ) 35.78 / 87.10 81.87 / 92.75 65.27 / 88.98 72.37 / 81.90 63.82 / 87.68 80.81 / 51.33 89.58 / 51.33 75.33 / 52.28 81.08 / 54.75 81.70 / 52.66
PRN (zhang2023prototypical, ) 79.26 / 84.30 80.41 / 87.60 76.93 / 77.02 75.87 / 76.53 78.12 / 81.36 80.90 / 76.48 88.84 / 84.79 67.82 / 51.70 74.63 / 70.57 78.05 / 70.89
Lisa (lai2023lisa, ) 87.37 / 84.31 93.34 / 91.73 84.45 / 85.28 85.51 / 87.72 87.67 / 87.26 88.66 / 79.18 83.57 / 72.87 57.77 / 35.73 77.75 / 69.01 76.94 / 64.20
AnomalyGPT (gu2023anomalygpt, ) 91.37 / 93.68 92.64 / 94.39 89.58 / 92.58 86.23 / 89.66 89.96 / 92.58 86.39 / 83.47 89.74 / 86.02 80.32 / 60.32 84.35 / 80.73 85.20 / 77.64
FabGPT (ours) 94.28 / 97.03 94.43 / 97.30 90.32 / 95.70 88.19 / 92.40 91.81 / 95.61 90.01 / 85.69 92.28 / 91.58 83.00 / 79.19 87.40 / 86.72 88.17 / 85.80
@@
Our experiment is conducted on the in-house SEM-WaD dataset. The dataset comprises images with a resolution of $480\times 480$, each accompanied by a corresponding mask image and related textual descriptions and analyses. We divide the data into training and test sets in a 7:3 ratio, both containing good and defective images.
@@
These baselines are also tuned on our SEM-WaD dataset, following their publicly available implementations and models. Among these baselines, Lisa (lai2023lisa, ) and AnomalyGPT (gu2023anomalygpt, ) are LMMs that support Q&A while the others do not.
@@ 4.2. Quantitative Results
Table 1 report the Image-AUC, Pixel-AUC, PRO, and AP values for different methods across 4 categories within the SEM-WaD dataset. It can be observed that our model outperforms all other methods in four evaluation metrics for most defect categories. For example, compared to AnomalyGPT (gu2023anomalygpt, ), which also employs prompt learning for fine-tuning, our model achieves a higher Image-AUC by 1.85% and a higher Pixel-AUC by 3.03%. Compared to the traditional detection model PRN (zhang2023prototypical, ), our model surpasses it by 10.12% in PRO and 14.91% in AP.
Table 2. The accuracies (%) of the models’ responses to various questions. “-” denotes the question is unsupported.
Questions Lisa (lai2023lisa, ) AnomalyGPT (gu2023anomalygpt, ) GPT-4 (achiam2023gpt, ) FabGPT
Presence 95.00 95.00 100.00 100.00
Category 92.50 75.00 37.50 97.50
Location - 72.50 87.50 95.00
Quantity - 77.50 80.00 95.00
Description 72.50 82.50 85.00 95.00
Analysis - - 90.00 97.50
Unrelated 20.00 12.00 98.00 98.00
All 70.00 69.08 82.57 96.86
Figure 5. Comparisons with non-LMM baselines.
Figure 6. Comparisons with LMM-based baselines.
@@ 4.3. Qualitative Results
According to Figure 7, FabGPT provides detailed defect knowledge, offering engineers valuable insights. Notably, the questions shown are not part of the training corpus. FabGPT identifies and localizes defects in SEM images, analyzing properties like ”hole.” It identifies root causes and suggests process solutions to optimize the production line. Additionally, FabGPT accurately answers general IC questions unrelated to the images, confirming the effectiveness of our modulation module in addressing modality bias.
@@ 4.4. Ablation Studies
Figure 7. Dialogue example of FabGPT for “hole” type defects. Questions with $\ast$” are general IC questions that are not closely related to the input images.
@@
Table 4 shows the advantages of our PM strategy. Cosine similarity outperforms matrix multiplication, and combining it with Linear and ReLU improves computation.
@@
Table 5 shows that our prompt embedding scheme effectively mitigates modality bias. By updating the corresponding coefficient “$a$” for visual tokens, our model effectively discerns the relationship between user queries and visual inputs and mitigates the modality bias issues.
@@ 5. Conclusions
In this paper, we introduce a novel large multimodal language model, FabGPT, for defect knowledge querying in the IC field, including defection, analysis, Q&A, etc. It employs three stages to gradually achieve the functionality of defect detection and high-quality dialogue. We validate the effectiveness of FabGPT on the SEM-WaD dataset and 100 questions. Our work provides great convenience for the semiconductor industry and also offers insights for further LMM research.
Acknowledgments
This work is sponsored by The National Natural Science Foundation of China (Grant No. 62034007).